OpenStudy (anonymous):

Algebra help!! What is the rate of change of the equation 5x+3y=-2? Explain how you found it.

OpenStudy (anonymous):

whats up

OpenStudy (anonymous):

um

OpenStudy (luckycoins888):

stuff

OpenStudy (anonymous):

er

OpenStudy (anonymous):

Let's solve for x. 5x+3y=−2 Step 1: Add -3y to both sides. 5x+3y+−3y=−2+−3y 5x=−3y−2 Step 2: Divide both sides by 5. 5x5=−3y−25 x=−35y+−25

OpenStudy (anonymous):

5x + 3y = -2 3y = -5x - 2 y = -(5/3)x - 2/3 The rate of change will always be the number infront of the x value. Therefore in this case the rate of change would be -5/3

OpenStudy (anonymous):

@felavin please explain more about how you got your answer. Like step by step

OpenStudy (anonymous):

do you know about rate?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

rate means slop? ok?

OpenStudy (anonymous):

Okay